OpenStudy (anonymous):

STEP BY STEP PLEASE: Divide: a. 4x = 32

OpenStudy (anonymous):

4x = 32 4x / 4 = 32 / 4 (divide both sides by 4) x = 8

OpenStudy (anonymous):

\[x=\frac{32}{4}=8\]